# Emulation Capacity between Idempotent Channels

### v1

@@ 4.1 Achievability in Theorem 1
This shows that $P_{\iota}=\iota(\mathds{1}_{\mathcal{A}})$ is an orthogonal projector in $\mathcal{B}$.
For $\widetilde{\iota}$, note that, as $\iota$ is subunital, $P_{\iota}\leq\mathds{1}_{\mathcal{B}}$, so that $\mathds{1}_{\mathcal{B}}-P_{\iota}\geq 0$ and $x\mapsto\iota(x)+\frac{\operatorname{\operatorname{tr}}(x)}{\operatorname{\operatorname{tr}}(\mathds{1}_{\mathcal{A}})}(\mathds{1}_{\mathcal{B}}-P_{\iota})$ is completely positive. The unitality is trivial.
Then, as $\iota^{-1}$ is a *-homomorphism, it is completely positive. Furthermore, as $P_{\iota}$ is an orthogonal projector, $x\in\mathcal{B}\mapsto P_{\iota}xP_{\iota}$ is completely positive. Therefore the composition $x\mapsto\iota^{-1}(P_{\iota}xP_{\iota})$ is completely positive as well. Then $\widetilde{\iota}^{-1}(\mathds{1}_{\mathcal{B}})=\iota^{-1}(P_{\iota})=\iota^{-1}(\iota(\mathds{1}_{\mathcal{A}}))=\mathds{1}_{\mathcal{A}}$.
Finally, for $x\in\mathcal{A}$, we have
$\widetilde{\iota}^{-1}\circ\widetilde{\iota}(x)$ $=\iota^{-1}(P_{\iota}\bigl(\iota(x)+\frac{\operatorname{\operatorname{tr}}(x)}{\operatorname{\operatorname{tr}}(\mathds{1}_{\mathcal{A}})}(\mathds{1}_{\mathcal{B}}-P_{\iota})\bigr)P_{\iota})$
$=\iota^{-1}(P_{\iota}\iota(x)P_{\iota})$
$=x,$
where we used for the last equality that $\operatorname{\operatorname{supp}}(\iota(x))\subseteq\operatorname{\operatorname{supp}}(P_{\iota})$ so that $P_{\iota}\iota(x)P_{\iota}=\iota(x)$. ∎
@@
with $\widetilde{\iota}$ and $\widetilde{\iota}^{-1}$ being defined as in Lemma 5. Indeed, for $x\in\mathcal{L}(\mathcal{H}_{\mathcal{F}}^{\otimes n})$,
$\widetilde{\iota}(\mathcal{F}^{*\otimes n}(x))=\iota(\mathcal{F}^{*\otimes n}(x))+\frac{\operatorname{\operatorname{tr}}(\mathcal{F}^{*\otimes n}(x))}{\operatorname{\operatorname{tr}}(\mathds{1}_{\mathcal{F}^{\otimes n}})}(\mathds{1}_{\mathcal{G}^{\otimes n}}-\iota(\mathds{1}_{\mathcal{F}^{\otimes n}}))\in\operatorname{Rg}(\mathcal{G}^{*\otimes n}),$
where $\mathds{1}_{\mathcal{F}^{\otimes n}}$ is the identity of $\mathcal{L}(\mathcal{H}_{\mathcal{F}}^{\otimes n})$ and $\mathds{1}_{\mathcal{G}^{\otimes n}}$ is the identity of $\mathcal{L}(\mathcal{H}_{\mathcal{G}}^{\otimes n})$. Since $\mathcal{G}^{*\otimes n}$ is idempotent, every element of its range is one of its fixed points, so that $\mathcal{G}^{*\otimes n}(\widetilde{\iota}(\mathcal{F}^{*\otimes n}(x)))=\widetilde{\iota}(\mathcal{F}^{*\otimes n}(x))$. We finally have by eq. (29) that $\widetilde{\iota}^{-1}\widetilde{\iota}(\mathcal{F}^{*\otimes n}(x))=\mathcal{F}^{*\otimes n}(x)$.
@@ 4.2 Converse of Theorem 1
Note that Theorems 4 and 6 are not converse one from another because the inequality between the $\ell_{p}$-norms of the shape vectors is strict in one case but not in the other.
@@
We write $\mathds{1}_{\mathcal{F}}$ and $\mathds{1}_{\mathcal{G}}$ the identity operator in $\mathcal{L}(\mathcal{H}_{\mathcal{F}})$ and $\mathcal{L}(\mathcal{H}_{\mathcal{G}})$, respectively. To simplify our notations, we also denote $e_{\mathcal{G}\mathcal{E}}=\mathcal{G}\mathcal{E}(\mathds{1}_{\mathcal{F}})^{0}$. Define
@@
Then, $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}$ restricted to the *-algebra $\operatorname{Rg}(\mathcal{F}^{*})$ is an injective unital *-homomorphism which embeds $\operatorname{Rg}(\mathcal{F}^{*})$ into $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$ as a subalgebra.
@@
Let $\Phi:\mathcal{L}(\mathcal{H})\rightarrow\mathcal{L}(\mathcal{H}^{\prime})$ be a quantum channel and write $e=\Phi(\mathds{1})^{0}$. Then
@@
    For all $x\in\mathcal{L}(\mathcal{H}^{\prime})$, $\Phi^{*}(x)=\Phi^{*}(exe)$.
@@
Let $y\in\mathcal{L}(\mathcal{H})$ be positive semidefinite. By positivity of $\Phi$,
$\Phi(y)\leq\|y\|\Phi(\mathds{1}).$
Therefore, $\operatorname{\operatorname{supp}}(\Phi(y))\subseteq\operatorname{\operatorname{supp}}(\Phi(\mathds{1}))=e$. Then, for all $x\in\mathcal{L}(\mathcal{H}^{\prime})$,
$\operatorname{\operatorname{tr}}(y\Phi^{*}(x))=\operatorname{\operatorname{tr}}(\Phi(y)x)=\operatorname{\operatorname{tr}}(e\Phi(y)ex)=\operatorname{\operatorname{tr}}(y\Phi^{*}(exe)).$
By linearity, this equation extends to all $y\in\mathcal{L}(\mathcal{H})$. Thus, $\Phi^{*}(x)=\Phi^{*}(exe)$.
@@
In particular, if $\mathcal{F}^{*}$ is subunital completely positive [^3] In fact, it suffices for $\mathcal{F}^{*}$ to be subunital $2$-positive for it to satisfy the Kadison-Schwarz inequality, see Section 5.2 of [18] for more details., it satisfies the Kadison-Schwarz inequality (eq. (34)) and we can use this second characterisation of its multiplicative domain.
@@
with $\widetilde{\mathcal{G}}^{*}$ defined in the statement of Proposition 6. Let $x\in\operatorname{Rg}(\mathcal{F}^{*})$. As $\operatorname{Rg}(\mathcal{F}^{*})$ is a *-algebra, $x^{*}x\in\operatorname{Rg}(\mathcal{F}^{*})$, and as $\mathcal{F}^{*}$ is idempotent, by Lemma 3, we have:
@@
Therefore, $x$ is in the multiplicative domain of $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}$. Thus, $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}$ restricted to $\operatorname{Rg}(\mathcal{F}^{*})$ is an injective unital *-homomorphism into the *-algebra $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$. It is injective as, for all $x\in\operatorname{Rg}(\mathcal{F}^{*})$,
$x=\mathcal{F}^{*}(x)=\mathcal{E}^{*}\mathcal{G}^{*}\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x).$
It is unital as it maps the identity of $\mathcal{L}(\mathcal{H}_{\mathcal{F}})$ to $e_{\mathcal{G}\mathcal{E}}$ which is the identity operator of the *-algebra $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$. Indeed, as both $\mathcal{G}$ and $\mathcal{D}$ are channels, $\mathcal{G}^{*}$ and $\mathcal{D}^{*}$ are unital so that $\mathcal{G}^{*}\mathcal{D}^{*}(\mathds{1}_{\mathcal{F}})=\mathds{1}_{\mathcal{G}}$ and $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(\mathds{1}_{\mathcal{F}})=e_{\mathcal{G}\mathcal{E}}$.
Now, to show that for all $p\in[1,+\infty]$, $\|\lambda(e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}})\|_{p}\leq\|\lambda(\operatorname{Rg}(\mathcal{G}^{*}))\|_{p}$, it suffices to remark that, by Proposition 1, we can decompose the Hilbert space $\mathcal{H}_{\mathcal{G}}$ as
@@
Hence, for all $k\in[K]$, the $k$-th coordinate of the shape vector of $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$ satisfies
@@
Then by Proposition 6, there is an injective unital *-homomorphism from $\operatorname{Rg}(\mathcal{F}^{*\otimes k})$ into the *-algebra $e_{\mathcal{G}^{\otimes n}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*\otimes n})e_{\mathcal{G}^{\otimes n}\mathcal{E}}$. Then, by Theorem 6 and the second statement of Proposition 6, we have for all $p\in[1,+\infty]$:
$\|\lambda(\mathcal{F}^{\otimes k})\|_{p}=\|\lambda(\operatorname{Rg}(\mathcal{F}^{*\otimes k}))\|_{p}\leq\|\lambda(e_{\mathcal{G}^{\otimes n}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*\otimes n})e_{\mathcal{G}^{\otimes n}\mathcal{E}})\|_{p}\leq\|\lambda(\operatorname{Rg}(\mathcal{G}^{*\otimes n}))\|_{p}.$
@@
$\|\lambda(\mathcal{F})\|_{p}^{k}\leq\|\lambda(\mathcal{G})\|^{n}_{p}.$
@@
5 Channel emulation with errors
@@
$\operatorname{\operatorname{tr}}(\mu\Phi_{1}\otimes\mathrm{Id}_{d}(\sigma))-\operatorname{\operatorname{tr}}(\mu\Phi_{2}\otimes\mathrm{Id}_{d}(\sigma))\leq\frac{1}{2}\|\Phi_{1}-\Phi_{2}\|_{\diamond}.$ (36)
@@
We begin with the case $p=1$. As we saw in Proposition 1, we can decompose $\mathcal{H}_{\mathcal{F}}$ as a direct sum
$\mathcal{H}_{\mathcal{F}}=\bigoplus_{k=1}^{K}\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2}$
@@
From Propososition 1, we can decompose $\mathcal{H}_{\mathcal{G}}$ and $\mathcal{G}$ as
@@
where $(a)$ follows from the fact that $\mathcal{D}\mathcal{G}_{2}:\mathcal{L}(\mathcal{H}_{\mathcal{G}_{1}})\rightarrow\mathcal{L}(\mathcal{H}_{\mathcal{F}})$ is a channel, so that $\mathcal{G}_{2}^{*}\mathcal{D}^{*}$ is unital. Then $(b)$ follows from the fact that for all $k\in[K]$ and $\nu\in[d_{k}]$, $|\nu\rangle\!\langle\nu|_{k}\otimes\frac{\mathds{1}_{m_{k}}}{m_{k}}$ is a state so that $\operatorname{\operatorname{tr}}(\mathcal{G}_{1}\mathcal{E}_{1}(|\nu\rangle\!\langle\nu|_{k}\otimes\frac{\mathds{1}_{m_{k}}}{m_{k}}))=1$, hence $\mathcal{G}_{1}\mathcal{E}_{1}(|\nu\rangle\!\langle\nu|_{k}\otimes\frac{\mathds{1}_{m_{k}}}{m_{k}})\leq\mathds{1}_{\mathcal{G}_{1}}$, and that as $\mathcal{G}_{2}^{*}\mathcal{D}^{*}$ is completely positive, $\mathcal{G}_{2}^{*}\mathcal{D}^{*}(|\nu\rangle\!\langle\nu|_{k}\otimes\mathds{1}_{m_{k}})\geq 0$. We then simply use that if $A,B\geq 0$ with $B\leq\mathds{1}$, $\operatorname{\operatorname{tr}}(A)\geq\operatorname{\operatorname{tr}}(AB)$. When injecting the last equation into eq. (41), we find
@@
We now move on to the proof of eq. (38), i.e. the case $p=+\infty$. To deal with the case where the blocks of the range of $\mathcal{F}^{*}$ would have multiplicities, we factorise $\mathcal{F}=\mathcal{F}_{2}\circ\mathcal{F}_{1}$ as we did for $\mathcal{G}$ in eq. (42), so that we have the decompositions
@@
$=\frac{1}{d^{2}}\operatorname{\operatorname{tr}}(\sum_{\nu,\eta,\nu^{\prime},\eta^{\prime}=1}^{d}|\nu\nu\rangle\!\langle\eta\eta||\nu^{\prime}\nu^{\prime}\rangle\!\langle\eta^{\prime}\eta^{\prime}|)$
@@
Let $\mathcal{F}$, $\mathcal{G}$ be two idempotent channels, $(k_{\nu})_{\nu\in\mathbb{N}}$, $(n_{\nu})_{\nu\in\mathbb{N}}$ be two integer sequences and $\operatorname{\varepsilon}>0$ satisfying the assumptions of Theorem 2. Let $p_{\min}\in\{1,+\infty\}$ such that

### v2

@@ 4.1 Achievability in Theorem 1
Finally, note that $\mathds{1}_{\mathcal{B}}-P_{\iota}$ is the orthogonal projector on $\ker(P_{\iota})$ which is orthogonal to $\operatorname{\operatorname{supp}}(P_{\iota})=\operatorname{Rg}(V)$, so that $(\mathds{1}_{\mathcal{B}}-P_{\iota})V=0$. Hence, for $x\in\mathcal{A}$,
$\widetilde{\iota}^{-1}(\widetilde{\iota}(x))=\widehat{\iota}^{-1}(E_{\widehat{\iota}(\mathcal{A})}(V^{*}\iota(E_{\mathcal{A}}(x))V))=\widehat{\iota}^{-1}(E_{\widehat{\iota}(\mathcal{A})}(\widehat{\iota}(x)))=\widehat{\iota}^{-1}(\widehat{\iota}(x))=x.$
@@
with $\widetilde{\iota}$ and $\widetilde{\iota}^{-1}$ being defined as in Lemma 6. Indeed, for $x\in\mathcal{L}(\mathcal{H}_{\mathcal{F}}^{\otimes n})$, $\widetilde{\iota}(\mathcal{F}^{*\otimes n}(x))\in\operatorname{Rg}(\mathcal{G}^{*\otimes n})$. Then, since $\mathcal{G}^{*\otimes n}$ is idempotent, every element of its range is one of its fixed points, so that $\mathcal{G}^{*\otimes n}(\widetilde{\iota}(\mathcal{F}^{*\otimes n}(x)))=\widetilde{\iota}(\mathcal{F}^{*\otimes n}(x))$. As $\mathcal{F}^{*\otimes n}(x)\in\operatorname{Rg}(\mathcal{F}^{*\otimes n})$, we can finally apply eq. (30) to get that $\widetilde{\iota}^{-1}\widetilde{\iota}(\mathcal{F}^{*\otimes n}(x))=\mathcal{F}^{*\otimes n}(x)$.
@@ 4.2 Converse of Theorem 1
Note that Theorems 5 and 7 are not converse one from another because the inequality between the $\ell_{p}$-norms of the shape vectors is strict in one case but not in the other.
@@
We write $\mathds{1}_{\mathcal{F}}$ and $\mathds{1}_{\mathcal{G}}$ the identity operator in $\mathcal{L}(\mathcal{H}_{\mathcal{F}})$ and $\mathcal{L}(\mathcal{H}_{\mathcal{G}})$ respectively. To simplify our notations, we also denote $e_{\mathcal{G}\mathcal{E}}=\mathcal{G}\mathcal{E}(\mathds{1}_{\mathcal{F}})^{0}$. Define
@@
Then, $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$ is a unital *-subalgebra of $\mathcal{L}(\mathcal{H}_{\mathcal{G}})$ whose identity operator is $e_{\mathcal{G}\mathcal{E}}$ and $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}$ restricted to the *-algebra $\operatorname{Rg}(\mathcal{F}^{*})$ is an injective unital *-homomorphism which embeds $\operatorname{Rg}(\mathcal{F}^{*})$ into $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$ as a subalgebra.
@@
Let $\Phi:\mathcal{L}(\mathcal{H})\rightarrow\mathcal{L}(\mathcal{H}^{\prime})$ be a quantum channel and write $e=\Phi(\mathds{1})^{0}$. Then,
@@
    for all $x\in\mathcal{L}(\mathcal{H}^{\prime})$, $\Phi^{*}(x)=\Phi^{*}(exe)$,
@@
Let $\mathcal{H}_{0}^{\prime}=\operatorname{\operatorname{supp}}(\Phi(\mathds{1}))$ and $V:\mathcal{H}_{0}^{\prime}\rightarrow\mathcal{H}^{\prime}$ be the isometry such that $e=VV^{*}$. Let $y\in\mathcal{L}(\mathcal{H})$ and $x\in\mathcal{L}(\mathcal{H}^{\prime})$, then
$\operatorname{\operatorname{tr}}(y\Phi^{*}(x))=\operatorname{\operatorname{tr}}(\Phi(y)x)=\operatorname{\operatorname{tr}}(VV^{*}\Phi(y)VV^{*}x)=\operatorname{\operatorname{tr}}(e\Phi(y)ex)=\operatorname{\operatorname{tr}}(y\Phi^{*}(exe)),$
where the second equality follows by Lemma 3. Therefore $\Phi^{*}(x)=\Phi(exe)$.
@@
In particular, if $\mathcal{F}^{*}$ is unital completely positive [^5] In fact, it suffices for $\mathcal{F}^{*}$ to be subunital $2$-positive for it to satisfy the Kadison-Schwarz inequality, see Section 5.2 of [19] for more details., it satisfies the Kadison-Schwarz inequality (eq. (35)) and we can use this second characterisation of its multiplicative domain.
@@
with $\widetilde{\mathcal{G}}^{*}$ defined in the statement of Proposition 7. Let $x\in\operatorname{Rg}(\mathcal{F}^{*})$. By the assumption that $\mathcal{F}=\widehat{\mathcal{F}}$, $\operatorname{Rg}(\mathcal{F}^{*})$ is a *-subalgebra of $\mathcal{L}(\mathcal{H}_{\mathcal{F}})$, thus $x^{*}x\in\operatorname{Rg}(\mathcal{F}^{*})$, and as $\mathcal{F}^{*}$ is idempotent, $xx^{*}$ is a fixed point of $\mathcal{F}^{*}$, so that we have:
@@
Therefore, $x$ is in the multiplicative domain of $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}$. Thus, $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}$ restricted to $\operatorname{Rg}(\mathcal{F}^{*})$ is an injective unital *-homomorphism into the *-algebra $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$, which is a *-subalgebra of $\mathcal{L}(\mathcal{H}_{\mathcal{G}})$ and whose identity operator is $e_{\mathcal{G}\mathcal{E}}$. It is injective as, for all $x\in\operatorname{Rg}(\mathcal{F}^{*})$,
$x=\mathcal{F}^{*}(x)=\mathcal{E}^{*}\mathcal{G}^{*}\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x),$
hence $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(x)=0$ implies $x=0$. It is unital as it maps the identity of $\mathcal{L}(\mathcal{H}_{\mathcal{F}})$ to $e_{\mathcal{G}\mathcal{E}}$ which is the identity operator of the *-algebra $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$. Indeed, as both $\mathcal{G}$ and $\mathcal{D}$ are channels, $\mathcal{G}^{*}$ and $\mathcal{D}^{*}$ are unital so that $\mathcal{G}^{*}\mathcal{D}^{*}(\mathds{1}_{\mathcal{F}})=\mathds{1}_{\mathcal{G}}$ and $\widetilde{\mathcal{G}}^{*}\mathcal{D}^{*}(\mathds{1}_{\mathcal{F}})=e_{\mathcal{G}\mathcal{E}}$.
Now, to show that $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$ is a *-subalgebra of $\mathcal{L}(\mathcal{H}_{\mathcal{G}})$ which satisfies for all $p\in[1,+\infty]$, $\|\lambda(e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}})\|_{p}\leq\|\lambda(\operatorname{Rg}(\mathcal{G}^{*}))\|_{p}$}, it suffices to remark that, by Proposition 1, we can decompose the Hilbert space $\mathcal{H}_{\mathcal{G}}$ as
@@
Hence, $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$ is a *-subalgebra for all $k\in[K]$, the $k$-th coordinate of the shape vector of $e_{\mathcal{G}\mathcal{E}}\operatorname{Rg}(\mathcal{G}^{*})e_{\mathcal{G}\mathcal{E}}$ satisfies
@@
By Proposition 3, then, if we denote $\widehat{\mathcal{F}^{\otimes k}}$ and $\widehat{\mathcal{G}^{\otimes n}}$ the reduced channels associated to $\mathcal{F}^{\otimes k}$ and $\mathcal{G}^{\otimes n}$, this condition is equivalent to
$\widehat{\mathcal{F}^{\otimes n}}=\mathcal{D}^{\prime}\widehat{\mathcal{G}^{\otimes n}}\mathcal{E}^{\prime},$
for some channels $\mathcal{D}^{\prime}$, $\mathcal{E}^{\prime}$. Thus, by Proposition 7, there is an injective unital *-homomorphism from $\operatorname{Rg}(\widehat{\mathcal{F}^{\otimes k}}^{*})$ into the *-algebra $e_{\widehat{\mathcal{G}^{\otimes n}}\mathcal{E}^{\prime}}\operatorname{Rg}(\widehat{\mathcal{G}^{\otimes n}}^{*})e_{\widehat{\mathcal{G}^{\otimes n}}\mathcal{E}^{\prime}}$, which is itself a *-subalgebra of $\mathcal{L}(\mathcal{H}_{\mathcal{G}})$. Then, by Theorem 7 and the second statement of Proposition 7, we have for all $p\in[1,+\infty]$:
$\|\lambda(\widehat{\mathcal{F}^{\otimes k}})\|_{p}=\|\lambda(\operatorname{Rg}(\widehat{\mathcal{F}^{\otimes k}}^{*}))\|_{p}\leq\|\lambda(e_{\widehat{\mathcal{G}^{\otimes n}}\mathcal{E}^{\prime}}\operatorname{Rg}(\widehat{\mathcal{G}^{\otimes n}}^{*})e_{\widehat{\mathcal{G}^{\otimes n}}\mathcal{E}^{\prime}})\|_{p}\leq\|\lambda(\operatorname{Rg}(\widehat{\mathcal{G}^{\otimes n}}^{*}))\|_{p}=\|\lambda(\widehat{\mathcal{G}^{\otimes n}}^{*})\|.$
@@
$\|\lambda(\mathcal{F})\|_{p}^{k}=\|\lambda(\widehat{\mathcal{F}})\|_{p}^{k}\leq\|\lambda(\widehat{\mathcal{G}})\|_{p}^{k}=\|\lambda(\mathcal{G})\|^{n}_{p}.$
@@
5 Emulation of idempotent channels with errors
@@
$\frac{1}{2}\|\Phi_{1}-\Phi_{2}\|_{\diamond}\geq\operatorname{\operatorname{tr}}(\mu\Phi_{1}\otimes\mathrm{Id}_{d}(\sigma))-\operatorname{\operatorname{tr}}(\mu\Phi_{2}\otimes\mathrm{Id}_{d}(\sigma)).$ (37)
@@
$\inf_{\mathcal{D},\mathcal{E}}\|\mathcal{F}-\mathcal{D}\mathcal{G}\mathcal{E}\|_{\diamond}=\inf_{\mathcal{D},\mathcal{E}}\|\widehat{\mathcal{F}}-\mathcal{D}\widehat{\mathcal{G}}\mathcal{E}\|_{\diamond}$.
Therefore, to simplify our notations, we will suppose without loss of generality that $\mathcal{F}$ and $\mathcal{G}$ are equal to their reduced version, i.e. $\mathcal{F}=\widehat{\mathcal{F}}$ and $\mathcal{G}=\widehat{\mathcal{G}}$. We prove separately the two following equations:
@@
We begin with the case $p=1$. As we saw in Proposition 1, we can decompose $\mathcal{H}_{\mathcal{F}}$ as an orthogonal direct sum
$\mathcal{H}_{\mathcal{F}}=\bigoplus_{k=1}^{K}\mathcal{H}_{k,1}\otimes\mathcal{H}_{k,2},$
@@
From Proposition 1, we can decompose $\mathcal{H}_{\mathcal{G}}$ and $\mathcal{G}$ as
@@
where $(a)$ follows from the fact that $\mathcal{D}\mathcal{G}_{2}:\mathcal{L}(\mathcal{H}_{\mathcal{G}_{1}})\rightarrow\mathcal{L}(\mathcal{H}_{\mathcal{F}})$ is a channel, so that $\mathcal{G}_{2}^{*}\mathcal{D}^{*}$ is unital. Then $(b)$ follows from the fact that for all $k\in[K]$ and $\nu\in[d_{k}]$, $|\nu\rangle\!\langle\nu|_{k}\otimes\frac{\mathds{1}_{m_{k}}}{m_{k}}$ is a state so that $\operatorname{\operatorname{tr}}(\mathcal{G}_{1}\mathcal{E}_{1}(|\nu\rangle\!\langle\nu|_{k}\otimes\frac{\mathds{1}_{m_{k}}}{m_{k}}))=1$, hence $0\leq\mathcal{G}_{1}\mathcal{E}_{1}(|\nu\rangle\!\langle\nu|_{k}\otimes\frac{\mathds{1}_{m_{k}}}{m_{k}})\leq\mathds{1}_{\mathcal{G}_{1}}$, and that as $\mathcal{G}_{2}^{*}\mathcal{D}^{*}$ is completely positive, $\mathcal{G}_{2}^{*}\mathcal{D}^{*}(|\nu\rangle\!\langle\nu|_{k}\otimes\mathds{1}_{m_{k}})\geq 0$. We then simply use that if $A\geq 0$ and $B\leq\mathds{1}$, $\operatorname{\operatorname{tr}}(A)\geq\operatorname{\operatorname{tr}}(AB)$. When injecting the last equation into eq. (42), we find
@@
We now move on to the proof of eq. (39), i.e. the case $p=+\infty$. To deal with the case where the blocks of the range of $\mathcal{F}^{*}$ have multiplicities, we factorise $\mathcal{F}=\mathcal{F}_{2}\circ\mathcal{F}_{1}$ as we did for $\mathcal{G}$ in eq. (43), so that we have the decompositions
@@
$=\frac{1}{d^{2}}\operatorname{\operatorname{tr}}(\sum_{\nu,\eta,\nu^{\prime},\eta^{\prime}=1}^{d}|\nu\nu\rangle\!\langle\eta\eta|\nu^{\prime}\nu^{\prime}\rangle\!\langle\eta^{\prime}\eta^{\prime}|)$
@@
Let $\mathcal{F}$, $\mathcal{G}$ be two idempotent channels, $(k_{\nu})_{\nu\in\mathbb{N}}$, $(n_{\nu})_{\nu\in\mathbb{N}}$ be two integer sequences and $\operatorname{\varepsilon}>0$ satisfying the assumptions of Theorem 2. Let $p_{\min}\in\{1,+\infty\}$ be such that